OpenStudy (anonymous):

More problems ugh! (Just 3)

OpenStudy (anonymous):

What is the value of the expression \[\frac{ 3 }{ 3^{-2} }\]

OpenStudy (anonymous):

@imasofty

Nnesha (nnesha):

\[\huge\rm \frac{ x^m }{ x^{-n} }=x^{m+n}\] exponent rule

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\[\frac{1}{b^{-n}}=b^n\] therefore \[\huge\frac{3}{3^{-2}}=3\times 3^2=3\times 3\times 3\]

OpenStudy (anonymous):

so b.27

Nnesha (nnesha):

if there are same bases at the numerator and 't the denominator you should move the denominator exponent to the top